OpenStudy (hocuspucus292333393):

A mixture contains 4 fluid ounces of water for every 3 fluid ounces of blue food coloring. How much blue food coloring needs to be added to 12 fluid ounces of water? 3 fluid ounces 9 fluid ounces 12 fluid ounces 16 fluid ounces

OpenStudy (hocuspucus292333393):

@jackthegreatest

OpenStudy (jackthegreatest):

ok it is a ratio

OpenStudy (jackthegreatest):

3/4 = x/12

OpenStudy (jackthegreatest):

@HocusPucus292333393

OpenStudy (hocuspucus292333393):

I thinks I was wrong about 4

OpenStudy (jackthegreatest):

whoopsie, i meant 12/4 which is 3 3 x 3 = 9

OpenStudy (hocuspucus292333393):

oh ha ha ok xD

OpenStudy (hocuspucus292333393):

that make since

OpenStudy (hocuspucus292333393):

so 9 ounces

OpenStudy (jackthegreatest):

yep

OpenStudy (hocuspucus292333393):

ok I got two more questions and that's all